### v2

# Quantum geometric tensor and wavepacket dynamics in two-dimensional non-Hermitian systems

Abstract
The quantum geometric tensor (QGT) characterizes the local geometry of quantum states, and its components directly account for the dynamical effects observed, e.g., in condensed matter systems. In this work, we address the problem of extending the QGT formalism to non-Hermitian systems with gain and loss. In particular, we investigate a wave-packet dynamics in two-band non-Hermitian systems to elucidate how non-Hermiticity affects the definition of QGT. We employ first-order perturbation theory to account for non-adiabatic corrections due to interband mixing. Our results suggest that two different generalizations of the QGT, one defined using only the right eigenstates and the other one using both the left and right eigenstates, both play a significant role in wave-packet dynamics. We then determine the accuracy of the perturbative approach by simulating a wave-packet dynamics in a well studied physical non-Hermitian system – exciton polaritons in a semiconductor microcavity. Our work aids deeper understanding of quantum geometry and dynamical behaviour in non-Hermitian systems.
preprint: APS/123-QED
Introduction.— The complex-valued quantum geometric tensor (QGT) was first proposed by Provost and Vallee in 1980 to quantify the distance (in parameter space) between two quantum states [1]. Its real (symmetric) part is the quantum metric tensor (QMT) describing the overlap between two quantum states and its imaginary (anti-symmetric) part is the Berry curvature describing the topology of the system’s energy bands. In periodic systems, e.g. solid-state lattices, the integral of the Berry curvature over a Brillouine zone is a topological invariant known as the Chern number [1, 2, 3, 4, 5, 6, 7, 8, 9]. The Berry curvature acts on the wave-packet centre of mass (e.g., an electron) as an effective magnetic field in momentum space, and results in an anomalous velocity transverse to the direction of the applied force [10, 11, 12, 13, 14, 15, 16, 17]. The QMT appears in the wave-packet equation of motions as perturbative corrections to the Berry curvature and describes the field induced positional shift [12, 13] and non-adiabatic effects [14]. It also plays an important role in flat-band superconductivity [18, 19, 20] and can also be assopciated with fidelity susceptibility that has been used to detect criticality, which plays an important role in quantum information theory [21, 22, 23, 24, 25].
Generalizing the QGT to systems effectively described by non-Hermitian Hamiltonians  [26, 27, 28, 29, 30, 31, 32, 33] is of fundamental importance for understanding the dynamical effects in systems with gain and loss. Non-Hermitian systems have attracted a great amount of interests because they can feature novel topological invariants [34], topological phases [34, 35], and topological edge states [36, 37, 38, 39, 40, 41, 42, 43, 44]. Non-Hermitian effects have also been experimentally studied in a variety of physical systems including cold atoms [45, 46], photonics [47, 48, 49, 50, 51], mechanical systems [52], electric circuits [53], magnons [54, 55], spin liquid [56], and exciton polaritons [57, 58].
The challenge in generalizing the QGT lies in the fact that non-Hermitian Hamiltonians have distinct left and right eigenstates [30, 31, 32, 33]. Some works therefore define the QGT using only the right eigenstates, which we denote the right-right (RR) formalism [59, 60, 61, 62, 63, 64], while others use both the left and right eigenstates, which we denote the left-right (LR) formalism [65, 66, 67, 68, 69, 70]. Previous studies proposed that it is the RR version for the Berry curvature that describes the anomalous velocity in the dynamics of a wave-packet centre of mass [71, 72, 73]. However, there are also non-Hermitian corrections arising from the Berry connections defined in both the RR and the LR formalism [71, 72, 73]. At the same time, while Refs. [59, 64] suggests that the RR QMT plays a significant role in the wave-packet dynamics, Ref. [70] states that the RR QMT is ill-defined due to the non-orthogonality of the right eigenstates. So far, it remains unclear which version of the QMT correctly represents the physical (measurable) quantities of non-Hermitian systems and how well the QMT describes the dynamical effects in non-Hermitian systems.
In this work, we address this ambiguity in the QGT definition by investigating the wave-packet dynamics of a two-band non-Hermitian system using the first-order perturbative theory.
We derive a semi-classical equation of the center-of-mass motion of a wave-packet, which explicitly shows the contributions of both the LR and RR components of the QGT. We then compared the analytical results with a numerical simulation using a model describing an experimentally accessible non-Hermitian system. The model describes two-dimensional hybrid light-matter particles in a solid state – microcavity exciton polaritons [74]. Both real and imaginary parts of the eigenenergies [58] as well as QGT components  [15, 75] can be measured in this system, in principle allowing for experimental verification of theoretical predictions.
Theory.— For an $n$-th level quantum state $|u_{n}\rangle$ that is smoothly dependent on momentum, the QGT can be defined as [1]:
$Q_{n,ij}=\langle\partial_{k_{i}}u_{n}|\partial_{k_{j}}u_{n}\rangle-\langle\partial_{k_{i}}u_{n}|u_{n}\rangle\langle u_{n}|\partial_{k_{j}}u_{n}\rangle$ (1)
where $k_{i},k_{j}$ denote the indices in momentum space. The QGT can also be written in terms of inter-band Berry connections $\mathbf{A}_{nm}=\langle u_{n}|i\partial_{\mathbf{k}}u_{m}\rangle$, in the form of $Q_{n,ij}=\sum_{m\neq n}A_{nm,i}A_{mn,j}$ [14, 76]. The QMT, $g_{n,ij}=\operatorname{Re}[Q_{n,ij}]$ and the Berry curvature, $\Omega_{n,ij}=-2\operatorname{Im}[Q_{n,ij}]$ correspond to the real and imaginary parts of the QGT, respectively.
In non-Hermitian systems, where the Hamiltonian $H\neq H^{\dagger}$, the left and right eigenstates defined as:
$H|u_{n}^{R}\rangle=E_{n}|u_{n}^{R}\rangle,\qquad\langle u_{n}^{L}|H=\langle u_{n}^{L}|E_{n}$ (2)
are not equal but are biorthogonal, satisfying $\langle u_{n}^{L}|u_{m}^{R}\rangle=0$ for $m\neq n$ [32, 31]. This distinction between the left and right eigenstates leads to different ways of defining the QGT. The QGT can be defined using only the right eigenstates, like in Refs. [59, 60, 61, 62, 63]. In this formalism, the inner product is defined as $\langle\cdot\rangle=\langle u_{n}^{R}|\cdot|u_{n}^{R}\rangle$ and the QGT and its components can be written as:
$\begin{split}Q_{n,ij}^{RR}=&\langle\partial_{k_{i}}u_{n}^{R}|\partial_{k_{j}}u_{n}^{R}\rangle-\langle\partial_{k_{i}}u_{n}^{R}|u_{n}^{R}\rangle\langle u_{n}^{R}|\partial_{k_{j}}u_{n}^{R}\rangle,\\ g_{n,ij}^{RR}=&\operatorname{Re}[Q_{n,ij}^{RR}],\qquad\Omega_{n,ij}^{RR}=-2\operatorname{Im}[Q_{n,ij}^{RR}].\end{split}$ (3)
In what follows, we will denote the components of the QGT defined this way as the right-right (RR) generalization. Similarly to the Hermitian case, the RR QMT and the RR Berry curvature correspond to the real and imaginary parts of the RR QGT, and both quantities are real-valued.
Alternatively, the QGT can also be defined using both the left and the right eigenstates, like in Refs. [66, 67, 68, 69, 70], where the inner product is defined using biorthogonality, $\langle\cdot\rangle=\langle u_{n}^{L}|\cdot|u_{n}^{R}\rangle$. The corresponding QGT and its components are
$\begin{split}Q_{n,ij}^{LR}=&\langle\partial_{k_{i}}u_{n}^{L}|\partial_{k_{j}}u_{n}^{R}\rangle-\langle\partial_{k_{i}}u_{n}^{L}|u_{n}^{R}\rangle\langle u_{n}^{L}|\partial_{k_{j}}u_{n}^{R}\rangle,\\ g_{n,ij}^{LR}=&\frac{1}{2}(Q_{n,ij}^{LR}+Q_{n,ji}^{LR}),\\ \Omega_{n,ij}^{LR}=&i(Q_{n,ij}^{LR}-Q_{n,ji}^{LR}).\end{split}$ (4)
We refer to this formalism as the left-right (LR) generalization. Unlike the Hermitian case, the LR QMT and the LR Berry curvature can be complex-valued, and correspond to the symmetric and the anti-symmetric parts of the LR QGT. Note that this is not the only way to define a LR generalization of the QMT. Ref. [70] proposed two other possible generalizations such as defining the LR QMT as the real part of the LR QGT or the real and symmetric part of the LR QGT. Here, we use Eqs. (4) to maintain the symmetric properties in the Hermitian case and our results support this definition, as will be shown below. Note that, in both formalisms, we define the eigenstates such that $\langle u_{n}^{L}|u_{n}^{R}\rangle=\langle u_{n}^{R}|u_{n}^{R}\rangle=1$ to ensure that both $Q_{n}^{RR}$ and $Q_{n}^{LR}$ are gauge-invariant.
The relevance of the components of the QGT to the dynamics has been investigated in both Hermitian and non-Hermitian systems. In the Hermitian case, the semi-classical equations of motion of a narrow wave-packet in a single band can be written as  [12, 13, 14]:
$\begin{split}\dot{\mathbf{k}}_{c}&=\mathbf{F}/\hbar\\ \dot{\mathbf{r}}_{c}&=\partial_{\mathbf{k}}\epsilon_{0}-\mathbf{F}\times\Big{(}\mathbf{\Omega}_{0}+\nabla_{\mathbf{k}}\times\frac{2g_{0}\cdot\mathbf{F}}{\epsilon_{1}-\epsilon_{0}}\Big{)},\end{split}$ (5)
where $\mathbf{k}_{c}$ and $\mathbf{r}_{c}$ are the center-of-mass (COM) momentum and position, respectively, and $\mathbf{F}$ is an external force. The first term, the gradient of the energy $\epsilon_{0}$ of the band, is the group velocity. The second term is a field-induced correction due to the (vector) Berry curvature $\mathbf{\Omega}_{0}$, resulting in the anomalous Hall drift [15, 14, 9, 77, 76] perpendicular to the field $\mathbf{F}$. The last term is the result of first-order perturbation theory accounting for non-adiabatic corrections due to a neighboring band with energy $\epsilon_{1}$. This term is second order in $\mathbf{F}$ and is proportional to the QMT $g_{0}$.
In the non-Hermitian case, only the first-order field-induced corrections were previously worked out [71, 72, 73] with the COM position written as:
$\dot{\mathbf{r}}_{c}=\partial_{\mathbf{k}}\operatorname{Re}\Big{[}\epsilon_{0}+\mathbf{F}\cdot(\mathbf{A}_{00}^{RR}-\mathbf{A}_{00}^{LR})\Big{]}-\mathbf{F}\times\mathbf{\Omega}_{0}^{RR}.$ (6)
Here, the real part of the complex-valued eigen-energy $\epsilon_{0}$ determines the group velocity and the RR Berry curvature determines the anomalous Hall drift. The non-Hermitian contribution is due to the real part of the difference of the RR and LR intra-band Berry connections of the band $A_{00}^{RR}$, $A_{00}^{LR}$ defined as
$\begin{split}\mathbf{A}_{nn}^{RR}=\langle u_{n}^{R}|i\partial_{\mathbf{k}}u_{n}^{R}\rangle,\qquad\mathbf{A}_{nn}^{LR}=\langle u_{n}^{L}|i\partial_{\mathbf{k}}u_{n}^{R}\rangle.\end{split}$ (7)
Note that $\mathbf{A}_{nn}^{LR}$ is generally complex-valued, but only the real part is relevant in Eq. (6). The Berry connection $\mathbf{A}_{nn}^{RR}$ is real-valued when the right eigenstates are normalized as $\langle u_{n}^{R}|u_{n}^{R}\rangle=1$ [71, 72, 75].
The non-Hermitian anomalous Berry connection [71, 73] $(\mathbf{A}_{nn}^{RR}-\mathbf{A}_{nn}^{LR})$ results in a field-induced correction to the energy[72]. In the Hermitian limit, this term in the equation of motion vanishes and $\mathbf{\Omega}^{RR}\rightarrow\mathbf{\Omega}$, reducing Eq. (6) to the Hermitian case [10], Eq. 5, without the adiabatic correction.
In this work, we derive the second-order field-induced corrections to find out how the non-Hermitian QGT components affect the wave-packet dynamics by including interband mixing.
Results.— We consider a perturbation to a two-dimensional (2D) Hamiltonian in the form of
$\tilde{H}=H-\mathbf{F}\cdot\mathbf{r},$ (8)
where $\mathbf{F}$ represents an external force and $\mathbf{r}$ is the position operator. We then consider a Gaussian wave packet in the state
$|W\rangle=\int w(\mathbf{k},t)e^{i\mathbf{k}\cdot\mathbf{r}}|\tilde{u}_{0}^{R}(\mathbf{k})\rangle d^{2}\mathbf{k},$ (9)
where $|\tilde{u}_{0}^{R}\rangle$ is a right eigenstate of the Hamiltonian $\tilde{H}$ that includes the external force.
In this work, we focus on two-band systems with gain and loss and denote $|\tilde{u}_{0}^{R}\rangle$ to be the eigenstate with the larger growth (or the smaller decay) rate. The wave packet will eventually evolve into the eigenstate with the largest $\operatorname{Im}E_{n}$ due to the loss or gain [72, 78]. Therefore, in order to employ a single-band approximation, we need to ensure that the wave packet stays in the eigenstate that has the larger $\operatorname{Im}E_{n}$, i.e. larger gain (or smaller loss).
We then apply the perturbation theory  [79] by expanding the eigenstate to the first order correction in $\mathbf{F}$ as,
$|\tilde{u}_{0}^{R}(\mathbf{k})\rangle=|u_{0}^{R}(\mathbf{k})\rangle-\frac{\mathbf{F}\cdot\mathbf{A}_{10}^{LR}}{\epsilon_{0}-\epsilon_{1}}|u_{1}^{R}(\mathbf{k})\rangle$ (10)
where $|u_{1}^{R}(\mathbf{k})\rangle$ is the other eigenstate (with larger loss), $\mathbf{A}_{10}^{LR}=\langle u_{1}^{L}|i\partial_{\mathbf{k}}u_{0}^{R}\rangle$ is the LR inter-band Berry connection, and $\epsilon_{0,1}$ are the complex-valued energy eigenvalues of the unperturbed Hamiltonian $H$. We emphasize that the inter-band Berry connection in the perturbative correction has to be defined in the “left-right”, instead of the conventional “right-right”, formalism due to biorthogonality in non-Hermitian systems [79].
Assuming that the wave packet is narrowly centered at its center of mass in momentum space $|w(\mathbf{k},t)|^{2}\approx\delta(\mathbf{k}-\mathbf{k}_{c})$, we then follow the formalism presented in Refs. [71, 73] and derive the equation of motion of the wave packet COM in position space up to the first-order perturbative correction (see SI for details) as:
$\begin{split}\hbar\dot{\mathbf{r}}_{c}=&\partial_{\mathbf{k}}\operatorname{Re}\Big{[}\epsilon_{0}+\mathbf{F}\cdot(\tilde{\mathbf{A}}_{00}^{RR}-\tilde{\mathbf{A}}_{00}^{LR})\Big{]}-\mathbf{F}\times\tilde{\mathbf{\Omega}}_{0}^{RR}\end{split}$ (11)
where,
$\tilde{\mathbf{A}}_{00}^{RR}=$ $\mathbf{A}_{00}^{RR}+\operatorname{Re}\bigg{[}\frac{2Q_{0}^{RR}\cdot\mathbf{F}}{\epsilon_{1}-\epsilon_{0}}\bigg{]}$
$-\operatorname{Im}\bigg{[}\partial_{\mathbf{k}}\Big{(}\frac{\mathbf{F}\cdot(\mathbf{A}_{00}^{RR}-\mathbf{A}_{00}^{LR})}{\epsilon_{1}-\epsilon_{0}}\Big{)}\bigg{]}$ (12)
$\tilde{\mathbf{A}}_{00}^{LR}=$ $\mathbf{A}_{00}^{LR}+\frac{2g^{LR}\cdot\mathbf{F}}{\epsilon_{1}-\epsilon_{0}}$ (13)
$\tilde{\mathbf{\Omega}}_{0}^{RR}=$ $\mathbf{\Omega}_{0}^{RR}+\nabla_{\mathbf{k}}\times\operatorname{Re}\bigg{[}\frac{2Q_{0}^{RR}\cdot\mathbf{F}}{\epsilon_{1}-\epsilon_{0}}\bigg{]}.$ (14)
These expressions form the main result of our work: Eq. (11) is the non-Hermitian generalization of Eq. (5) and Eqs. (12, 13, 14), explicitly showing the first-order perturbation corrections to Eq. (6). In the following, we will show that the second-order (in $\mathbf{F}$) corrections have contributions from the QMT, Berry curvature, and anomalous non-Hermitian Berry connection.
Let us first examine the correction to the Berry curvature $\tilde{\mathbf{\Omega}}_{0}^{RR}$ by comparing Eq. (11) to Eq. (5). The second-order field-induced correction clearly transforms as $g_{0}/(\epsilon_{1}-\epsilon_{0})\rightarrow\operatorname{Re}[Q_{0}^{RR}/(\epsilon_{1}-\epsilon_{0})]$. Note, however, that this is not simply $g_{0}\rightarrow g_{0}^{RR}$ due to the complex denominator. In fact, there is an additional non-Hermitian correction which comes from the RR Berry curvature itself, as shown below:
$\operatorname{Re}\Big{[}\frac{2Q_{0}^{RR}\cdot\mathbf{F}}{\Delta\epsilon}\Big{]}=\operatorname{Re}\left[\frac{1}{\Delta\epsilon}\right]2g_{0}^{RR}\cdot\mathbf{F}-\operatorname{Im}\left[\frac{1}{\Delta\epsilon}\right]\mathbf{F}\times\mathbf{\Omega}_{0}^{RR}$ (15)
This new term depends on the imaginary part of the complex energy gap, which naturally vanishes in the Hermitian limit or along the imaginary Fermi arc, where $\operatorname{Im}\Delta\epsilon=0$ [58, 78].
The second-order field-induced corrections to the anomalous non-Hermitian Berry connection have two components, i.e. $\operatorname{Re}[\mathbf{F}\cdot(\tilde{\mathbf{A}}_{00}^{RR}-\tilde{\mathbf{A}}_{00}^{LR})]=I_{0}+I_{1}+I_{2}$, where $I_{0}=\mathbf{F}\cdot(\mathbf{A}_{00}^{RR}-\mathbf{A}_{00}^{LR})]$ is the first-order term. The first correction is proportional to the difference of the RR and LR QMTs and has the form
$I_{1}=2\mathbf{F}\cdot\operatorname{Re}\left[\frac{g_{0}^{RR}-g_{0}^{LR}}{\Delta\epsilon}\right]\cdot\mathbf{F},$ (16)
while the second term is proportional to the anomalous Berry connection and has the form
$I_{2}=-\mathbf{F}\cdot\operatorname{Im}\bigg{[}\partial_{\mathbf{k}}\Big{(}\frac{\mathbf{F}\cdot(\mathbf{A}_{00}^{RR}-\mathbf{A}_{00}^{LR})}{\Delta\epsilon}\Big{)}\bigg{]}.$ (17)
Note that both the real and imaginary parts of the LR QMT and Berry connections contribute to the corrections due to the complex denominator (see SI for the explicit forms of $I_{1}$ and $I_{2}$). This novel non-Hermitian effect shows that the LR QMT should indeed be defined as a complex-valued quantity [66, 67, 70]. This is in contrast to the corrections to the Berry curvature, Eq. (14), which originate from the real-valued RR QMT and Berry curvature.
To examine the corrections closely, we integrate Eq. 11 to get (see SI for the derivation):
$\mathbf{r}_{c}=$ $\bigg{[}-\partial_{\mathbf{k}}\varphi+\tilde{\mathbf{A}}_{00}^{RR}\bigg{]}_{\mathbf{k}=\mathbf{k}_{c}},$ (18)
$\varphi=$ $-\frac{1}{\hbar}\int_{t_{0}}^{t}\operatorname{Re}\bigg{[}\epsilon_{0}-\mathbf{F}\cdot\tilde{\mathbf{A}}_{00}^{LR}\bigg{]}_{\mathbf{k}=\mathbf{k}_{c}}dt^{\prime}.$ (19)
The positional shift [12] of the COM, given by Eq. (18), is due to the gradient of the phase $\varphi$ and the corrected RR Berry connection. The latter, which is given by Eq. (12), has two field-induced corrections. The first depends on the RR QGT, $Q_{0}^{RR}$, which is the same term that appears in field-induced correction to the Berry curvature $\tilde{\mathbf{\Omega}}_{0}^{RR}$ in Eq. (14). The second correction depends on the anomalous non-Hermitian Berry connection, $\mathbf{A}_{00}^{RR}-\mathbf{A}_{00}^{LR}$, which also appears in the field-induced correction to the energy in Eq. (6). There is also a field-induced correction to the geometric (Berry) phase [2, 71, 73], Eq. (19), which is proportional to the LR QMT, $g_{0}^{LR}$.
In the Hermitian limit, the eigenenergies become real-valued, $\mathbf{A}_{00}^{LR}=\mathbf{A}_{00}^{RR}$, and $Q_{0}^{RR}=Q_{0}^{LR}$ becoming Hermitian as well. Therefore, the correction to the LR and RR Berry connections both take the forms of $\mathbf{A}_{00}\rightarrow\mathbf{A}_{00}+2(g\cdot\mathbf{F})/(\Delta\epsilon)$, and we recover the previous results given by Eq. (5[12, 13, 14].
Figure 1: The (a) real and (b) imaginary parts of the eigenenergies of the exciton polaritons in a gapped phase with $\Delta_{z}=0.007$ eV. (c, d) The cross-sections of the real and imaginary parts of the eigenenergies along the $k_{y}$-axis, respectively. The orange line and surface represent the upper band, the blue line and surface represent the lower band, and the black dots represent the initial momentum of the wave-packet center-of-mass. (e, f) The RR quantum metric and Berry curvature. (g, h) The real and imaginary parts of the LR quantum metric. The blue and red lines in (e, g, h) show the $xy$-components and the $yy$-components respectively, since these are the only components that appears in the dynamics when the force is in the $y$-direction.
To demonstrate the relevance of the QGT components to the description of dynamical effects in a two-band non-Hermitian system, we simulate the wave-packet dynamics in an exciton-polariton system formed by strong coupling of an optically active, 2D-confined exciton to a photon confined in an optical microcavity. It is a 2D (two-dimensional) system featuring rich non-Hermitian effects due to inherent loss and gain [80]. The non-Hermiticity of the system arises from the decay of polaritons, mainly due to photons leaking out through the imperfect mirrors of the microcavity. Due to a pseudospin (polarisation) degree of freedom, this system can also feature non-trivial bands with exceptional points in momentum space [58, 78, 75]. In principle, exciton-polaritons can provide an experimental platform for verifying the results presented here since components of both the QGT [9, 15] and its non-Hermitian generalization [75] can be measured in this system.
We employ split-step methods [81, 82] to verify that the first order perturbation theory developed here improves the accuracy of the semi-classical equation of motion. We consider a two-band non-Hermitian effective Hamiltonian for exciton polaritons in a form of a complex-valued effective magnetic field $\overrightarrow{\mathbf{G}}=[G_{x},G_{y},G_{z}]$ acting on the exciton-polariton pseudospins described by the $2\times 2$ Pauli matrices $\overrightarrow{\sigma}=[\sigma_{x},\sigma_{y},\sigma_{z}]$. The Hamiltonian has the form [58, 78]
$\begin{split}H(\mathbf{k})&=H_{0}(\mathbf{k})+\overrightarrow{\mathbf{G}}(\mathbf{k})\cdot\overrightarrow{\sigma}\\ H_{0}(\mathbf{k})&=E_{0}+\frac{\hbar^{2}k^{2}}{2m}-i\gamma_{0}-i\gamma_{2}k^{2}-i\gamma_{4}k^{4}\\ \overrightarrow{\mathbf{G}}(\mathbf{k})&=[\tilde{\alpha}+\tilde{\beta}(k_{x}^{2}-k_{y}^{2}),2k_{x}k_{y}\tilde{\beta},\Delta_{z}],\end{split}$ (20)
[…]
We initialized the wave-packet at $\mathbf{k}=(0,0.9)\mu$m-1 in the upper eigenstate which has the larger $\operatorname{Im}E$ (or less loss) [see the dot in Figs. 1(c,d)]. The two bands are separated at this point in momentum space and $\partial_{k_{x}}\operatorname{Re}\epsilon_{0}=0$. Experimentally, the wave-packet can be prepared using resonant optical excitation, where the exciting laser wavelength is adjusted to match the energy and the excitation angle is tuned to match the momentum of the target state [15].
An external force $\mathbf{F}=(0,5\times 10^{-4})$ eV/$\mu$m is applied accelerating the wave-packet along the $+k_{y}$-axis towards smaller (real) energy spacing. In an optical microcavity, this force naturally arises from the small variation or tilt in the cavity length introduced by the sample fabrication process, which translates to a linear energy gradient experienced by polaritons [85]. As the bands get closer, the mixing between the two eigenstates will be more profound and can result in oscillation due to the zitterbewegung effect [86, 77]. However, since the wave packet evolves into the less dissipative eigenstate, the inter-band mixing will gradually disappear and so will the oscillation. We also ensure that the wave packet does not cross the imaginary Fermi arc [58], where the imaginary parts of the eigenenergies cross, otherwise the single-band approximation will fail [78] as the loss rates of the eigenstates will switch. The numerical results are presented as dots in Figs. 2.
In this simple case, Eq. (11) becomes:
$v_{x}=$ $\frac{1}{\hbar}\partial_{k_{x}}\operatorname{Re}\Big{[}F_{y}\Big{(}\tilde{A}_{00,y}^{RR}-\tilde{A}_{00,y}^{LR}\Big{)}\Big{]}-\frac{F_{y}}{\hbar}\tilde{\Omega}_{0,xy}^{RR}$ (21)
$v_{y}=$ $\frac{1}{\hbar}\partial_{k_{y}}\operatorname{Re}\Big{[}\epsilon_{0}+F_{y}\Big{(}\tilde{A}_{00,y}^{RR}-\tilde{A}_{00,y}^{LR}\Big{)}\Big{]}.$ (22)
The explicit forms of $\tilde{A}_{00,y}^{RR}-\tilde{A}_{00,y}^{LR}$ and $\tilde{\Omega}_{0,xy}^{RR}$ are presented in the SI.
The solution of Eq. (11) gives a better fit to the numerical results than the solution of Eq. (6). This is apparent from the discrepancy between the numerical and analytical solutions plotted in Figs. 2 (c,d). Although the nonadiabatic correction is rather small in our example, a larger force or a smaller energy gap will result in a much larger correction, meaning that the previously derived equation of motion without perturbative corrections Eq. (6) will no longer be adequately describing the dynamics. In this case, however, we might also need to extend our treatment beyond the first-order perturbation theory and consider higher order corrections. On the other hand, if the wave packet propagates for a longer time, it would cross the imaginary Fermi arc, and the semi-classical equations of motion will no longer be valid [78].
Figure 2: (a, d) The real-space trajectory of exciton-polariton wave-packets initially placed in the upper eigenstate in 3 different gapped phases with $\Delta_{z}=0.007$ eV. Shown are the simulated results (black dots), the solution of semi-classical equation of motion derived previously Eq. (6) (blue lines) and the solution of the semi-classical equation of motion derived here Eq. (11) (purple lines). (c, d) The difference between the analytical results and the numerical simulation with the same color scheme as in (a, b).
Conclusion.— We derive the first-order perturbative corrections to the semi-classical equation of motion describing the centre-of-mass motion of a wave-packet of a non-Hermitian two-band system. Our results show that both the left-right (LR) and right-right (RR) quantum geometric tensor (QGT) play a role in the dynamics. In particular, the RR QGT describes the non-Hermitian generalization of the field-induced positional shift, while the LR QGT describes the field-induced correction to the Berry phase. Interestingly, our results also suggest novel second-order anomalous Hall drift correction arising from the RR Berry curvature, and the positional-shift correction arising from the difference between the RR and LR QMT and the anomalous Berry connection. More importantly, our theoretical results show that the imaginary part of the LR QMT and Berry connections play a direct role in the second-order field-induced corrections, unlike the first-order case. Our results settle the dispute on whether the LR or the RR QGT should be used to describe the dynamics of non-Hermitian systems, and instead suggest that both LR and RR QGTs represent physical quantities and play a significant role in the dynamics of non-Hermitian systems.
We further confirmed the accuracy of our results by simulating the center-of-mass dynamics of a non-Hermitian exciton-polariton system with a gapped spectrum, which can be realized in microcavity exciton-polaritons exposed to a strong out-of-plane magnetic field [87, 88] or hosted in liquid-crystal cavities [89].
Our work opens up many possibilities for further investigation in the QGT in non-Hermitian systems. In particular, our results should be relevant for the physics of second-harmonic generation [90], non-linear response theory [91, 92], quantum information [21, 22] and flat-band superconductivity [18, 19, 20], where the quantum metric tensor plays a significant role.
We acknowledge support from the Australian Research Council (ARC) through the Centre of Excellence Grant CE170100039 and the Discovery Early Career Researcher Award DE220100712, and Australian Government Research Training Program (RTP) Scholarship.